OpenStudy (anonymous):

(please help I got the answer wrong) A thief steals a number of rare plants from a nursery. On the way out, the thief meets three security guards, one after another. To each security guard, the thief is forced to give one-half of the plants that he still had, plus 2 more. Finally, the thief leaves the nursery with 11 plants . How many plants were originally stolen? (I got 116, pretty sure that's wrong)

OpenStudy (anonymous):

88

OpenStudy (anonymous):

since he lost half his plant 3 time you muli 11*2*2*2=88

OpenStudy (anonymous):

Not the correct answer but thank you for helping anyway!

OpenStudy (anonymous):

it might be 104

OpenStudy (anonymous):

or 90

OpenStudy (anonymous):

i think 116 is correct. He left with 11 Add 2 to get 13 Double is 26 Add 2 for 28 Double is 56 Add 2 is 58 Double is 116

OpenStudy (anonymous):

It was 116 thank you for clearing that out!

OpenStudy (anonymous):

74

OpenStudy (anonymous):

he left with 11 3rd guard taken 9 ( thief 2 more) before 3rd guard he had 11+9=20 2nd guard taken 18(thief 2 more) before 2nd guard 20+18=38 1st guard taken 36(thief 2 more) he stolen 36+38=74 does this correct